OpenStudy (anonymous):

Write 2 equations that are parallel to the line y=2/3x + 4

Directrix (directrix):

Parallel lines have the same slope but a different y-intercept. The given line y=(2/3) x + 4 is in y = mx + b form where m is the slope and b is the y-intercept. The slope of the given line is (2/3). That is the slope you want to use in the two linear equations you create. Choose any number as the y-intercept as long as you do not choose 4. @Educate_Me

OpenStudy (anonymous):

Okay , thanks .